An infinite line of charge with uniform line charge density of 1C/m is placed along the y -axis.A charge of 1C is placed on the x -axis at a distance of d=3m from the origin.At what distance (r) from the origin on the x -axis the total electric field is zero. (Assume 0<r<d)
ANSWER:2m
1
Expert's answer
2020-01-20T05:23:53-0500

The electric field of an infinite line charge is given by


"E_1(r)=\frac{\lambda}{2\pi r \varepsilon_0}"

The electric field of point charge placed at distance "d" from the origin is given by


"E_2(r)=\frac{q}{4\pi \varepsilon_0 (r-d)^2}"

At the point where the electric field is zero


"E_1(r)=E_2(r)"

Hence

"\frac{\lambda}{2\pi r \varepsilon_0}=\frac{q}{4\pi \varepsilon_0 (r-d)^2}"

After some algebra, we obtain an equation

"r^2-(2d+\frac{q}{2\lambda})r+d^2=0"

"r^2-6.5r+9=0"

Roots:


"r=2,\quad r=4.5"

For "0<r<d" we obtain "r=2\:\rm m."